# Spin kinetic theory with a nonlocal relaxation time approximation

### v1

@@ I Introduction
In the past years, relativistic kinetic theory with spin has been extensively studied Weickgenannt et al. (2019); Gao and Liang (2019); Hattori et al. (2019a); Wang et al. (2019); Weickgenannt et al. (2021a); Liu et al. (2020); Weickgenannt et al. (2021b); Sheng et al. (2021); Wagner et al. (2022, 2023). As already outlined in an old seminal paper for the nonrelativistic case Hess and Waldmann (1966) and recently demonstrated for the relativistic case in Refs. Weickgenannt et al. (2021a, b), where the nonlocal collision term for spin-1/2 particles was obtained from the Wigner-function formalism, it is essential to take into account the nonlocality of particle collisions to describe the conversion between orbital angular momentum and spin in kinetic theory: the nonlocal part of the collision term is responsible for aligning spin polarization along the vorticity direction, while the local part drives spin diffusion. [^1] Spin diffusion here refers to the process of eliminating a possible initial inhomogeneous polarization through the transport of spin-carrying particles, while the spin itself is conserved in particle collisions. Due to the complexity of solving kinetic theory directly, it is often convenient to study instead equations of motion for spin hydrodynamics Florkowski et al. (2018a, b, 2019c); Montenegro and Torrieri (2019); Hattori et al. (2019b); Bhadury et al. (2021); Singh et al. (2021); Montenegro and Torrieri (2020); Gallegos et al. (2021); Fukushima and Pu (2021); Li et al. (2021); Wang et al. (2021); Hu (2022); Hongo et al. (2021); Daher et al. (2022); Weickgenannt et al. (2022a, b); Gallegos et al. (2022); Cao et al. (2022); Weickgenannt et al. (2022c); Biswas et al. (2023); Weickgenannt (2023); Weickgenannt and Blaizot (2024); Daher et al. (2024); Wagner et al. (2024); Drogosz et al. (2024); Wagner (2024), which may be obtained, e.g., from the Boltzmann equation. In Ref. Weickgenannt et al. (2022a) second-order spin hydrodynamics was derived from nonlocal spin kinetic theory. It was found that the nonlocal collision term contributes new terms to the polarization vector in the Navier-Stokes limit, which are proportional to both the difference between the thermal vorticity and the spin potential, and to the thermal shear. However, a numerical implementation of the equations of motion of Ref. Weickgenannt et al. (2022a) is still missing and may be demanding due to the presence of various complicated collision integrals. Thus, for practical applications both of spin kinetic theory and hydrodynamics it may be useful to provide a simple approximation of the nonlocal collision term, which models its essential microscopic properties.
@@ II Spin kinetic theory with a nonlocal collision term
The dynamics of a system of massive spinning particles in the kinetic regime can be described by a distribution function $f(x,p,\mathfrak{s})$, which depends on the space-time $x$, the particle four-momentum $p$ and the particle spin four-vector $\mathfrak{s}$. Its evolution is determined by the on-shell Boltzmann equation Weickgenannt et al. (2021a, b)
@@
where $E_{p}\equiv\sqrt{\mathbf{p}^{2}+m^{2}}$ is the particle energy and $t\equiv(1,\mathbf{0})$ is the time unit vector. The exact, rather complicated, form of $\Delta^{\mu}$ can be found in Ref. Wagner et al. (2022). Note that $\Delta$ appears here naturally as being of order $\hbar$.
As discussed in Refs. Weickgenannt et al. (2021a, b, 2022a), the nonlocal collision term features specific properties related to the notion of equilibrium. In spinless kinetic theory, local equilibrium is defined as a state where the collision term vanishes, while the left-hand side of the Boltzmann equation (1) is nonzero. On the other hand, in global equilibrium both sides of the Boltzmann equation vanish separately. The same concepts apply to spin kinetic theory if the collision term is local, i.e., if all shifts $\Delta^{\mu}$ in Eq. (2) are set to zero. In that case, the local-equilibrium distribution function becomes a function of the sum of the collisional invariants [^2] We assume here that the fluid is uncharged. $p^{\mu}$ and $\Sigma_{\mathfrak{s}}^{\mu\nu}\equiv-(1/m)\epsilon^{\mu\nu\alpha\beta}p_{\alpha}\mathfrak{s}_{\beta}$, where the latter quantity is the so-called dipole-moment tensor which is conserved in local collisions, [^3] Note that the whole formalism presented in this paper is valid only for massive particles. each contracted with a Lagrange multiplier. Up to first order in $\hbar$, the local-equilibrium distribution function then reads
@@ III The nonlocal relaxation time approximation
which defines the spin potential. Note that $\Delta^{\mu}$ is of first order in $\hbar$, while $\Sigma_{\mathfrak{s}}^{\mu\nu}$ is of zeroth order, and $f$ and $f_{\text{LE}}$ contain terms both of zeroth and first order. However, zeroth-order terms in $f$ and $f_{\text{LE}}$ are independent of $\mathfrak{s}^{\mu}$, such that their contributions vanish when performing the spin integration, since the terms in the first round brackets are linear in $\mathfrak{s}^{\mu}$.
@@ V Asymptotic polarization in relativistic heavy-ion collisions
$\mathcal{A}^{\mu}_{\infty}=-\frac{\hbar}{4m}\left[\tilde{\Omega}^{\mu\nu}p_{\nu}+\frac{2}{E_{p}+m}\xi\epsilon^{\mu\lambda\rho\sigma}p_{\rho}{t}_{\sigma}p^{\nu}(\partial_{\lambda}\beta_{\nu}+\Omega_{\lambda\nu})\right]f^{(0)}_{\text{LE}}\;.$ (22)
Comparing with Eq. (21), we find agreement if we set $\xi=(E_{p}+m)/E_{p}$ and $\Omega^{\mu\nu}=\varpi^{\mu\nu}$ in Eq. (22).
@@
We see from comparison to Eq. (22) that $\zeta$ then is directly related to $\xi=(1-\zeta)(E_{p}+m)/E_{p}$. In particular, the long-time polarization (22) corresponds to the polarization obtained from Zubarev with the HW pseudo-gauge if $\xi=0$. Using the HW pseudo-gauge in the Zubarev formalism therefore corresponds to a situation in which the conversion between orbital angular momentum and spin happens very slowly, such that the nonlocal part of the collision term can be neglected and the spin tensor is considered to be conserved. Applying pseudo-gauge transformations with different values of $\zeta$ in Eq. (27), one can thus choose the equivalent of different time scales for angular-momentum conversion in the Zubarev formalism. We remark that this relation between the pseudo gauge and the angular-momentum conversion time scale is specific to the Zubarev formalism, where only the local-equilibrium part of the density matrix is considered and interaction terms are neglected in general.
@@ VI.3 Kinetic-theory solution for a rotating cylinder
with $\beta$ being constant, $\partial_{z}\mathbf{u}_{\bot}=0$, and $\partial^{(x}u^{y)}=0$. Formally, this distribution function corresponds to a local-equilibrium distribution function, which does not have contributions at first order in $\hbar$ since we choose the initial polarization to be zero. However, in general, it is not identical to the local-equilibrium distribution function (4), which contains a nonzero spin potential even at the initial time. In fact, $f_{\text{LE}}\neq f_{\text{in}}$ at any time, and in particular $f(t=0)\equiv f_{\text{in}}\neq f_{\text{LE}}(t=0)$. We will come back to the inital conditions later. We note that the nonlocality enters the collision term only at first order in $\hbar$, hence, the zeroth-order distribution function is not affected and retains its initial equilibrium form. We may therefore replace $f^{(0)}$ by $f^{(0)}_{\text{LE}}$ in Eq. (36),

### v2

@@ I Introduction
In the past years, relativistic kinetic theory with spin has been extensively studied Weickgenannt et al. (2019); Gao and Liang (2019); Hattori et al. (2019a); Wang et al. (2019); Weickgenannt et al. (2021a); Liu et al. (2020); Weickgenannt et al. (2021b); Sheng et al. (2021); Wagner et al. (2022, 2023); Fang and Pu (2024); Wang and Lin (2024). As already outlined in an old seminal paper for the nonrelativistic case Hess and Waldmann (1966) and recently demonstrated for the relativistic case in Refs. Weickgenannt et al. (2021a, b), where the nonlocal collision term for spin-1/2 particles was obtained from the Wigner-function formalism, it is essential to take into account the nonlocality of particle collisions to describe the conversion between orbital angular momentum and spin in kinetic theory: the nonlocal part of the collision term is responsible for aligning spin polarization along the vorticity direction, while the local part drives spin diffusion. [^1] Spin diffusion here refers to the process of eliminating a possible initial inhomogeneous polarization through the transport of spin-carrying particles, while the spin itself is conserved in particle collisions. Due to the complexity of solving kinetic theory directly, it is often convenient to study instead equations of motion for spin hydrodynamics Florkowski et al. (2018a, b, 2019c); Montenegro and Torrieri (2019); Hattori et al. (2019b); Bhadury et al. (2021); Singh et al. (2021); Montenegro and Torrieri (2020); Gallegos et al. (2021); Fukushima and Pu (2021); Li et al. (2021); Wang et al. (2021); Hu (2022); Hongo et al. (2021); Daher et al. (2022); Weickgenannt et al. (2022a, b); Gallegos et al. (2022); Cao et al. (2022); Weickgenannt et al. (2022c); Biswas et al. (2023); Weickgenannt (2023); Weickgenannt and Blaizot (2024); Daher et al. (2024); Wagner et al. (2024); Drogosz et al. (2024); Wagner (2024), which may be obtained, e.g., from the Boltzmann equation. In Ref. Weickgenannt et al. (2022a) second-order spin hydrodynamics was derived from nonlocal spin kinetic theory. It was found that the nonlocal collision term contributes new terms to the polarization vector in the Navier-Stokes limit, which are proportional to both the difference between the thermal vorticity and the spin potential, and to the thermal shear. However, a numerical implementation of the equations of motion of Ref. Weickgenannt et al. (2022a) is still missing and may be demanding due to the presence of various complicated collision integrals. Thus, for practical applications both of spin kinetic theory and hydrodynamics it may be useful to provide a simple approximation of the nonlocal collision term, which models its essential microscopic properties.
@@ II Spin kinetic theory with a nonlocal collision term
The dynamics of a system of massive spinning particles in the kinetic regime can be described by a distribution function $f(x,p,\mathfrak{s})$, which depends on the space-time $x$, the particle four-momentum $p$ and the particle spin four-vector $\mathfrak{s}$. This distribution function is defined in the extended phase space $(x,p,\mathfrak{s})$ and is directly related to components of the Wigner function. It was shown in Ref. Weickgenannt et al. (2021a) that the constraint equations for the Wigner function restrict the momentum to its mass-shell and the spin four vector to be orthogonal to momentum. Furthermore, the dynamical equations for the Wigner function yield the equation of motion for $f$. One finds that its evolution is determined by the on-shell Boltzmann equation Weickgenannt et al. (2021a, b)
@@
where $E_{p}\equiv\sqrt{\mathbf{p}^{2}+m^{2}}$ is the particle energy and $t\equiv(1,\mathbf{0})$ is the time unit vector. The exact, rather complicated, form of $\Delta^{\mu}$ can be found in Ref. Wagner et al. (2022). Note that $\Delta$ appears here naturally as being of order $\hbar$. We remark that for the derivation of Eq. (2) one employs standard assumptions of kinetic theory, in particular a low-density approximation.
As discussed in Refs. Weickgenannt et al. (2021a, b, 2022a), the nonlocal collision term features specific properties related to the notion of equilibrium. In spinless kinetic theory, local equilibrium is defined as a state where the collision term vanishes, while the left-hand side of the Boltzmann equation (1) is nonzero. On the other hand, in global equilibrium both sides of the Boltzmann equation vanish separately. The same concepts apply to spin kinetic theory if the collision term is local, i.e., if all shifts $\Delta^{\mu}$ in Eq. (2) are set to zero. In that case, the local-equilibrium distribution function becomes a function of the sum of the collisional invariants [^2] We assume here that the fluid is uncharged. $p^{\mu}$ and $\Sigma_{\mathfrak{s}}^{\mu\nu}\equiv-(1/m)\epsilon^{\mu\nu\alpha\beta}p_{\alpha}\mathfrak{s}_{\beta}$, where the latter quantity is the so-called dipole-moment tensor which is conserved in local collisions, each contracted with a Lagrange multiplier. Note that the whole formalism presented in this paper is valid only for massive particles. Up to first order in $\hbar$, the local-equilibrium distribution function then reads
@@ III The nonlocal relaxation time approximation
which defines the spin potential. Note that $\Delta^{\mu}$ is of first order in $\hbar$, while $\Sigma_{\mathfrak{s}}^{\mu\nu}$ is of zeroth order, and $f$ and $f_{\text{LE}}$ contain terms both of zeroth and first order. However, zeroth-order terms in $f$ and $f_{\text{LE}}$ are independent of $\mathfrak{s}^{\mu}$, such that their contributions vanish when performing the spin integration, since the terms in the first round brackets are linear in $\mathfrak{s}^{\mu}$. We remark that in principle, one could solve the Boltzmann equation with the NLRTA even without imposing any matching condition. However, in that case, the conservation laws would be violated and the Lagrange multipliers $\beta^{\mu}$ and $\Omega^{\mu\nu}$ would remain undetermined. Therefore, it is crucial to impose Eqs. (10) and (11).
@@ V Asymptotic polarization in relativistic heavy-ion collisions
$\mathcal{A}^{\mu}_{\infty}=-\frac{\hbar}{4m}\left[\tilde{\Omega}^{\mu\nu}p_{\nu}+\frac{2}{E_{p}+m}\xi\epsilon^{\mu\lambda\rho\sigma}p_{\rho}{t}_{\sigma}p^{\nu}(\partial_{\lambda}\beta_{\nu}+\Omega_{\lambda\nu})\right]f^{(0)}_{\text{LE}}\;,$ (22)
a relation which is independent of pseudo gauge. Comparing Eq. (22) with Eq. (21), we find agreement if we set $\xi=(E_{p}+m)/E_{p}$ and $\Omega^{\mu\nu}=\varpi^{\mu\nu}$ in Eq. (22).
@@
We see from comparison to Eq. (22) that $\zeta$ then is directly related to $\xi=(1-\zeta)(E_{p}+m)/E_{p}$. In particular, the long-time polarization (22) corresponds to the polarization obtained from Zubarev with the HW pseudo-gauge if $\xi=0$. Recalling that the polarization obtained from the Zubarev formalism is pseudo-gauge dependent, since an explicit form of the spin tensor is used, we conclude that the HW pseudo-gauge in the Zubarev formalism corresponds to a situation in which the conversion between orbital angular momentum and spin happens very slowly, such that the nonlocal part of the collision term can be neglected and the spin tensor is considered to be conserved. Applying pseudo-gauge transformations with different values of $\zeta$ in Eq. (27), one can thus choose the equivalent of different time scales for angular-momentum conversion in the Zubarev formalism. We remark that this relation between the pseudo gauge and the angular-momentum conversion time scale is specific to the Zubarev formalism, where only the local-equilibrium part of the density matrix is considered and interaction terms are neglected in general.
@@ VI.3 Kinetic-theory solution for a rotating cylinder
with $\beta$ being independent of spatial coordinates, $\partial_{z}\mathbf{u}_{\bot}=0$, and $\partial^{(x}u^{y)}=0$. Formally, this distribution function corresponds to a local-equilibrium distribution function, which does not have contributions at first order in $\hbar$ since we choose the initial polarization to be zero. However, in general, it is not identical to the local-equilibrium distribution function (4), which contains a nonzero spin potential even at the initial time. In fact, $f_{\text{LE}}\neq f_{\text{in}}$ at any time, and in particular $f(t=0)\equiv f_{\text{in}}\neq f_{\text{LE}}(t=0)$. We will come back to the inital conditions later. We note that the nonlocality enters the collision term only at first order in $\hbar$, hence, the zeroth-order distribution function is not affected and retains its initial equilibrium form. This implies that fluid velocity and temperature obtained from the matching conditions are constant with time. We may therefore replace $f^{(0)}$ by $f^{(0)}_{\text{LE}}$ in Eq. (36),